OpenStudy (anonymous):

Solve using matrices: x+y-z=-4 3x-y+z=0 -x+3y-2z=1 I NEED THE ANSWER... !!!!

OpenStudy (anonymous):

(1,-4,-7)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

do you know the gaussian elimination method?

OpenStudy (anonymous):

Yes, I was using it to solve your question.

OpenStudy (anonymous):

Ok, I have another one that I would like you to look at!!! :)

OpenStudy (anonymous):

Ok. sure and If you have confused about Gaussian elliminate. People perform on youtube are really good job.

OpenStudy (anonymous):

x+y-z=2 -x-4y+4z=-14 -x-3y-2z=15

OpenStudy (anonymous):

(-2,9,5)

OpenStudy (anonymous):

Thank you... If I put these into a gaussian elimination calculator... are the variables (the x'x and y's) do they represent a or a ?

OpenStudy (anonymous):

I worked on the paper, i dont know how to use the calculator, sorry!